# Learning Task-Agnostic Motifs to Capture the Continuous Nature of Animal Behavior

### v2

@@ 1 Introduction
A critical direction in animal behavior research has been identifying recurring patterns, often referred to as stereotyped behavioral syllables, like back grooming, running, and sniffing, directly from large-scale behavior recordings. Behavior segmentation methods (Wiltschko et al., 2015; Weinreb et al., 2024; Luxem et al., 2022; Hsu & Yttri, 2021; Berman et al., 2014) seek to uncover such structured patterns in behavior by dividing continuous pose trajectories into discrete syllables. Classic behavior segmentation approaches can be categorized into three groups: (1) supervised classification (Marks et al., 2022; Segalin et al., 2021), (2) clustering-based methods (Hsu & Yttri, 2021; Berman et al., 2014; Whiteway et al., 2021), and (3) switching-dynamics-based methods (Wiltschko et al., 2015; Weinreb et al., 2024; Luxem et al., 2022; Costacurta et al., 2022). The segmented syllables can then be used to build structured representations of movement for downstream neurobehavioral study.
@@
However, existing behavior segmentation methods face four major limitations. First, they rely on the assumption that behavior consists of discrete action syllables, which oversimplifies the inherently continuous nature of movement and introduces ambiguity during action transitions. Second, they often extract actions as abstract syllables that fail to capture how individual body parts contribute to different motions. For example, back and side grooming both involve similar forelimb movements combined with different turning dynamics, and sniffing may occur while walking or sitting, with similar head motion but distinct lower-body patterns. Capturing such compositional structure across body parts is essential for a more detailed understanding of natural behavior. Third, most segmentation models are either non-generative (e.g., clustering (Hsu & Yttri, 2021; Berman et al., 2014)) or rely on restrictive generative assumptions (e.g., linear dynamics and Markov models (Wiltschko et al., 2015; Weinreb et al., 2024; Luxem et al., 2022)), often leading to unrealistic synthesized behaviors. Fourth and most importantly, classification/clustering-based methods ignore temporal dependencies and fail to capture motifs reflecting true pose dynamics, while dynamics-based methods impose linear or nonlinear dynamics assumptions that may not match actual animal behavior, making such assumptions potentially harmful.
@@
To address these limitations, we introduce a new perspective: modeling behavior under the reinforcement learning (RL) framework. We study behavioral dynamics and motor motifs by inferring the animal’s policy through an RL-based imitation learning (IL) framework. It not only enables more realistic behavior generation through RL but also allows us to discover reusable motor motif sets to construct a policy driven by internal rewards. By viewing behavior through this lens, we gain a more flexible, generative, and interpretable understanding of motor motifs that go beyond the constraints of discrete segmentation. Note that Aldarondo et al. (2024) also applied RL-based IL to analyze animal behavior, but without parsing long untrimmed behaviors into fine-grained, interpretable motor motifs, so their work lies outside the scope of behavior segmentation considered here.
@@
We hypothesize that animals draw from a fixed set of core motor motifs to construct diverse movements over long behavioral trajectories (Santuz et al., 2019; Flash & Hochner, 2005). Building on this, we propose Motif-based Continuous Dynamics discovery (MCD) to parse long trajectories and uncover motifs and policies that reflect behavioral dynamics. Concretely, we learn interpretable latent representations, or motif sets, via spectral decomposition–based representation learning in RL (Dai et al., 2014; Ren et al., 2023; Shribak et al., 2024). These motifs correspond to low-level motion patterns serving as modular building blocks of behavior and can involve different body parts. For instance, face grooming (forepaw-to-face) and body grooming (torso strokes) share grooming motifs while engaging distinct body parts. These motifs can then be used to sufficiently represent policies that characterize complex high-level behaviors. Finally, we apply imitation learning to train motif-based policies from demonstrations. This framework leverages RL in two aspects: (1) motifs are inferred through RL-based representation learning, and (2) we use policies formed from motifs to characterize behavioral dynamics. As shown later, both aspects avoid any model assumptions while capturing motifs and policies that faithfully reflect behavioral dynamics.
@@
Another key innovation in constructing policies from motifs is that each motif’s contribution evolves continuously over time, reflecting dynamic behavioral changes. Some motifs may be brief, others prolonged, and multiple motifs can be active simultaneously to build ongoing movement (Fig. 1). This flexible, compositional view cannot be achieved with traditional discrete syllables. MCD thus provides a nuanced account of how complex actions arise from dynamic motif combinations and enables testing whether fine-grained motifs depend on specific neural circuits. Compared with prior segmentation methods, MCD offers soft segmentation that captures continuous time-varying processes rather than discrete switches.
Note that there is a rich literature in robotics on learning skills through imitation learning (Paraschos et al., 2013; Lioutikov et al., 2017; Li et al., 2017; Ajay et al., 2021; Peng et al., 2022; Kuang et al., 2025). We use RL-based imitation not to compete with existing skill learning methods, as most are not suited to the behavior segmentation task for neuroscience study in this paper. Rather, our goal is to employ the RL framework as a principled way to characterize the continuous nature of animal behavior while rendering fine-grained motor motifs, analogous to how Keypoint-MoSeq relies on an SLDS framework (Weinreb et al., 2024) and VAME relies on an autoencoder framework (Luxem et al., 2022). To summarize, we contribute to behavior understanding for neuroscience and neuroethology through the following points:
@@
  • Our method reveals the continuous nature of animal behavior, moving beyond the discrete segmentation assumptions of existing methods.
  • It provides a nuanced understanding of how complex behaviors emerge from dynamic combinations of fundamental motor motifs.
@@ 2 Preliminaries
$Q_{h}(s,a)=r(s,a)+\gamma\,\mathbb{E}_{s^{\prime}\sim P(\cdot\mid s,a)}\bigl[V_{h+1}(s^{\prime})\bigr],\quad V_{h}(s)=\mathbb{E}_{a\sim\pi(\cdot\mid s)}\bigl[Q_{h}(s,a)\bigr].$ (1)
@@ 3 Motif-based Continuous Dynamics (MCD) Discovery
Spectral decomposition has been widely studied in RL representation learning (Ren et al., 2023; Zhang et al., 2022; Shribak et al., 2024). We adopt this approach here to define motifs given the transition kernel and define rewards accordingly. Since spectral decomposition derives latents directly from the transition kernel without model assumptions, motif learning is thus independent of model assumptions and faithfully reflects the motifs present in the behavior data.
Given the motif definition, substituting Eq. 2 and the linearized reward model into the Bellman equation (Eq. 1), we get:
$Q(s,a)=r(s,a)+\gamma\!\!\int\!V(s^{\prime})P(s^{\prime}|s,a)\mathrm{d}s^{\prime}=\phi(s,a)^{\top}\!\!\left[w+\gamma\!\!\int\!V(s^{\prime})\mu(s^{\prime})q(s^{\prime})\mathrm{d}s^{\prime}\right]=\phi(s,a)^{\top}u,\!$ (3)
@@
$\pi(a|s)=\arg\max_{\pi}\;\;\left[\mathbb{E}_{\pi}[Q(s,a)]+H(\pi)\right]=\frac{\exp(\phi(s,a)^{\top}u)}{\sum_{a^{\prime}\in\mathcal{A}}\exp(\phi(s,a^{\prime})^{\top}u)},$ (4)
where $H(\pi):=\sum_{a\in\mathcal{A}}\pi(a|s)\,\log\bigl(\pi(a|s)\bigr)$ is the entropy.
Proposition 1.
@@
From Def. 1 and Prop. 1, we conclude that the learned motifs and policies do not rely on model assumptions, yet the policies faithfully capture behavioral dynamics as action tendencies conditioned on state. Thus, our method is assumption-free while capturing dynamics, unlike classification/clustering methods (no dynamics) or dynamics-based methods (restrictive assumptions).
Next, we introduce how to learn $\phi(s,a)$ and $\mu(s^{\prime})$ (motif discovery), as well as $u$ (motif-based policy learning that characterizes the continuous behavioral dynamics) from demonstrations. The learning procedure differs depending on the nature of the behavior data (i.e. discrete or continuous).
@@ 3.1 Discrete Version
Motif discovery. For discrete state-action spaces, we apply spectral methods such as singular value decomposition (SVD) (Golub & Reinsch, 1971; Golub & Van Loan, 2013; Trefethen & Bau, 2022) or spectral decomposition representation (Ren et al., 2023; HaoChen et al., 2021) to learn the representations $\phi(s,a),\mu(s^{\prime})=\arg\min_{\phi,\mu}||P(s^{\prime}|s,a)-\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime})||^{2}$. The resulting motif set $\phi(s,a)$ is then used in the subsequent policy learning stage.
@@
$\max_{u}\;\;\mathbb{E}_{(s,a)\sim\tau^{e}}\left[\log\pi(a|s)\right]=\max_{u}\;\;\mathbb{E}_{(s,a)\sim\tau^{e}}\left[\log\frac{\exp(\phi(s,a)^{\top}u)}{\sum_{a^{\prime}\in\mathcal{A}}\exp\left(\phi(s,a^{\prime})^{\top}u\right)}\right].$ (5)
@@ 3.2 Continuous Version
While learning from discrete data is relatively straightforward, the continuous case presents additional challenges for two main reasons. First, in the motif discovery step (Eq. 2), the decomposition $P(s^{\prime}|s,a)=\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime})$ is too restrictive to capture the complexity of continuous behavioral dynamics, such as pose transitions in freely moving animals (Weinreb et al., 2024). For example, consider a common and biologically plausible behavioral model: $s^{\prime}=h(s,a)+\epsilon$, where $h$ is a dynamics function and $\epsilon$ is Gaussian noise. This additive structure, widely used in behavioral modeling, contrasts with the multiplicative form $\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime})$, suggesting that parameterizations preserving additive relationships between $\{s,a\}$ and $s^{\prime}$ are more suitable. Second, in motif-based policy learning, for discrete datasets with a small action space, the denominator (partition function) in Eq. 5 is easy to compute. But for continuous data, the action space is infinite, making it infeasible to enumerate all actions and integrate. In light of these challenges, we adopt an alternative approach to learn the motif representations and policy in continuous state-action spaces.
@@
$P(s^{\prime}|s,a)=q(s^{\prime})\exp\left(\psi(s,a)^{\top}\nu(s^{\prime})-\log Z(s,a)\right),Z(s,a)=\int q(s^{\prime})\exp(\psi(s,a)^{\top}\nu(s^{\prime}))ds^{\prime},$ (6)
@@
Proposition 2 (Connection to Motif Definition).
@@
Given the EBM model in Eq. 6, the transition kernel can be approximated by $P(s^{\prime}|s,a)\approx\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime})$, where $\phi(s,a)\in\mathbb{R}^{d}$ is an explicit function of $\psi(s,a)$ and the partition function $Z(s,a)$, and $\mu(s^{\prime})\in\mathbb{R}^{d}$ is a function of $\nu(s^{\prime})$.
Appendix. C contains the full proof and derivation of $\phi$ and $\mu$ in terms of $\psi$, $\nu$, and $Z$.
@@
To learn $\psi$ and $\nu$, we employ noise-contrastive estimation (NCE) (Ma & Collins, 2018; Gutmann & Hyvärinen, 2010; 2012), which enables optimization of unnormalized statistical models without explicitly computing the partition function. In this way, we sidestep the intractable computation of $Z(s,a)$ in Eq. 6 by solving
@@
$\min_{\psi,\,\nu}\;\;\mathbb{E}_{\begin{subarray}{c}(s,a)\sim\tau^{e},s^{\prime}\sim P(\cdot|s,a),\\ s^{\prime\prime}_{i}\sim\rho\end{subarray}}\left[\log\frac{\exp(\psi(s,a)^{\top}\nu(s^{\prime}))}{\exp(\psi(s,a)^{\top}\nu(s^{\prime}))+\Sigma_{i=1}^{k}\exp(\psi(s,a)^{\top}\nu(s^{\prime\prime}_{i}))}\right],$ (7)
@@
where $s^{\prime}$ denotes a positive sample drawn from the transition distribution $P(\cdot|s,a)$, and $s^{\prime\prime}_{i}$ for $i=1,\ldots,k$ are negative samples from distribution $\rho$.
@@
Connection to behavioral dynamics. With simple algebra, we obtain the quadratic potential function $P(s^{\prime}|s,a)\propto q(s^{\prime})\exp\left(\|\psi(s,a)\|^{2}/2\right)\exp\left(-\|\psi(s,a)-\nu(s^{\prime})\|^{2}/2\right)\exp\left(\|\nu(s^{\prime})\|^{2}/2\right)$ from Eq. 6. By enforcing unit-norm constraints $\|\psi(s,a)\|^{2}=\|\nu(s^{\prime})\|^{2}=1$, assuming $Z(s,a)$ as a constant and $q(s^{\prime})$ as uniform distribution, as well as taking $\nu$ to be the identity map, we obtain a generalized Gaussian form: $P(s^{\prime}|s,a)=\frac{1}{Z}\exp\left(-\|s^{\prime}-\psi(s,a)\|^{2}\right)$, which aligns with commonly adopted assumptions in animal behavior modeling discussed earlier. Thus, Eq. 6 offers a more general framework that extends traditional dynamics models for studying behavior. Moreover, in practice, using Eq. 6 results in a unimodal distribution over $s^{\prime}$, which closely matches the empirical structure observed in $P(s^{\prime}|s,a)$ from animal behavior data. Thus, while Eq. 2 is theoretically valid in continuous domains, we adopt the parameterization in Eq. 6 for continuous state and action spaces, as it more effectively supports the learning of motif representations underlying animal behavior.
@@
Motif-based policy learning. After we obtain the representation $\psi(s,a)$ and $\nu(s^{\prime})$ from Eq. 7, we could theoretically get motif sets $\phi(s,a)$ expressed as the function of $\psi(s,a)$ and a normalizing term $Z(s,a)$ (see Appendix. C). However, since in practice $Z(s,a)$ remains intractable, even with the optimal $\psi$ it is still hard to obtain $\phi$ exactly. Therefore, we introduce a mapping $f:\psi\rightarrow\phi$, parameterized with a neural network, and learn it via policy learning. Our aim is to learn a function $f$ so that $\phi=f(\psi)$ yields optimal basis functions of policy that best account for the animal behavior data. By applying $\phi=f(\psi)$ to Eq. 3, we obtain $Q(s,a)=f\bigl(\psi(s,a)\bigr)^{\top}u$.
@@
As mentioned earlier, learning both $f$ and $u$ using the MLE objective in Eq. 5 becomes intractable for continuous data, as the denominator involves integration over an unbounded continuous action space. This brings us back to the challenge of estimating an unnormalized energy function, $\pi(a|s)\propto\exp(Q(s,a))\propto\exp(f(\psi(s,a))^{\top}u)$. Thus, it’s reasonable to apply NCE here again,
@@
$\min_{f,\,u}\;\;\mathbb{E}_{\begin{subarray}{c}(s,a)\sim\tau^{e},\\ (s_{i}^{\prime},a^{\prime}_{i})\sim\tau^{e}\end{subarray}}\left[\log\frac{\exp(f(\psi(s,a))^{\top}u)}{\exp(f(\psi(s,a))^{\top}u)+\Sigma_{i=1}^{k}\exp(f(\psi(s,a^{\prime}_{i}))^{\top}u)}\right],$ (8)
@@
where $\{s,a\}$ are positive samples and $\{s^{\prime}_{i},a^{\prime}_{i}\}$ are negative samples.
3.3 Understanding Animal Behavior via Motif and Policy Learning

### v3

@@ Abstract
behavioral neuroscience, animal behavior, compositional behavior modeling, behavior segmentation, neuroethology
@@ 1 Introduction
A critical direction in animal behavior research has been identifying recurring patterns, often referred to as stereotyped behavioral syllables, like back grooming, running, and sniffing, directly from large-scale behavior recordings. Behavior segmentation methods (Wiltschko et al., 2015; Weinreb et al., 2024; Luxem et al., 2022; Hsu and Yttri, 2021; Berman et al., 2014) seek to uncover such structured patterns in behavior by dividing continuous pose trajectories into discrete syllables. Classic behavior segmentation approaches can be categorized into three groups: (1) supervised classification (Marks et al., 2022; Segalin et al., 2021), (2) clustering-based methods (Hsu and Yttri, 2021; Berman et al., 2014; Whiteway et al., 2021), and (3) hidden-Markov-model(HMM) based methods (Wiltschko et al., 2015; Weinreb et al., 2024; Luxem et al., 2022; Costacurta et al., 2022). The segmented syllables can then be used to build structured representations of movement for downstream neurobehavioral study.
@@
However, existing behavior segmentation methods (Wiltschko et al., 2015; Weinreb et al., 2024; Luxem et al., 2022; Hsu and Yttri, 2021; Berman et al., 2014) overlook several features of the behavior data. First, continuity, they model continuous behavior as combinations of discrete action syllables, which oversimplifies the inherently continuous nature of movement and introduces ambiguity during action transitions. Therefore, they may fail to capture the details of delicate behavior dynamics. Second, compositionality, they often extract complex coordinated body movements as abstract syllables that fail to capture how individual body parts contribute to different motions. Therefore, they fail to reveal the connections and distinctions between syllables. For example, back and side grooming both involve similar forelimb movements combined with different turning dynamics, and sniffing may occur while walking or sitting, with similar head motion but distinct lower-body patterns. Third, long-term dependency. They either ignore temporal dependency between actions or only consider a very short time window when encoding the behavior. Therefore, they often fail to capture the long-term and multi-scale property of syllables (See Appendix K for discussions on temporal dependency). Apart from these explicit features, most models are either non-generative (e.g., clustering (Hsu and Yttri, 2021; Berman et al., 2014)) or rely on restrictive generative assumptions (e.g., linear dynamics and Markov models (Wiltschko et al., 2015; Weinreb et al., 2024; Luxem et al., 2022)), often leading to unrealistic synthesized behaviors.
@@
To address these limitations, we introduce a new perspective: modeling behavior under the reinforcement learning (RL) framework. We study behavioral dynamics and motor motifs by inferring the animal’s policy through an RL-based imitation learning (IL) framework. It not only enables more realistic behavior generation through RL but also allows us to discover reusable motor motif sets to construct a policy driven by internal rewards. By viewing behavior through this lens, we gain a more flexible, generative, and interpretable understanding of motor motifs that go beyond the constraints of discrete segmentation. Note that (Aldarondo et al., 2024) also applied RL-based IL to analyze animal behavior, but without parsing long untrimmed behaviors into fine-grained, interpretable motor motifs, so their work lies outside the scope of behavior segmentation considered here.
@@
We hypothesize that animals draw from a fixed set of core motor motifs to construct diverse movements over long behavioral trajectories (Santuz et al., 2019; Flash and Hochner, 2005). Building on this, we propose Motif-based Continuous Dynamics discovery (MCD) to parse long trajectories and uncover motifs and policies that reflect behavioral dynamics. Concretely, we learn interpretable latent representations, or motif sets, via spectral decomposition–based representation learning in RL (Dai et al., 2017; Ren et al., 2022; Shribak et al., 2024). These motifs correspond to low-level motion patterns serving as modular building blocks of behavior and can involve different body parts. For instance, face grooming (forepaw-to-face) and body grooming (torso strokes) share grooming motifs while engaging distinct body parts. These motifs can then be used to sufficiently represent policies that characterize complex high-level behaviors. Finally, we apply imitation learning to train motif-based policies from demonstrations. This framework leverages RL in two aspects: (1) motifs are inferred through RL-based representation learning, and (2) we use policies formed from motifs to characterize behavioral dynamics. As shown later, both aspects avoid any model assumptions while capturing motifs and policies that faithfully reflect behavioral dynamics.
@@
Another key innovation in constructing policies from motifs is that each motif’s contribution evolves continuously over time, reflecting dynamic behavioral changes. Some motifs may be brief, others prolonged, and multiple motifs can be active simultaneously to build ongoing movement (Fig. 1). This flexible, compositional view cannot be achieved with traditional discrete syllables. MCD thus provides a nuanced account of how complex actions arise from dynamic motif combinations and enables testing whether fine-grained motifs depend on specific neural circuits. Compared with prior segmentation methods, MCD offers soft segmentation that captures continuous time-varying processes rather than discrete switches. Further, theoretically the horizon in RL is infinite, enabling longer temporal dependency modeling.
Note that there is a rich literature in robotics on learning skills through reinforcement learning (Appendix I). We use RL-based imitation not to compete with existing skill learning methods, as most are not suited to the behavior segmentation task for neuroscience study in this paper. Rather, our goal is to employ the RL framework as a principled way to characterize the continuous nature of animal behavior while rendering fine-grained motor motifs, analogous to how Keypoint-MoSeq relies on an SLDS framework (Weinreb et al., 2024) and VAME relies on an autoencoder framework (Luxem et al., 2022). To summarize, we contribute to behavior understanding for neuroscience and neuroethology through the following points:
@@
  • Our method reveals the continuous, compositional and long-term dependent nature of animal behavior, providing a nuanced understanding of how complex behaviors emerge from dynamic combinations of fundamental motor motifs. It moves beyond the discrete segmentation assumptions of existing methods.
@@ 2 Preliminaries
$Q_{h}(s,a)$ $=r(s,a)+\gamma\,\mathbb{E}_{s^{\prime}\sim P(\cdot\mid s,a)}\bigl[V_{h+1}(s^{\prime})\bigr],$
$V_{h}(s)$ $=\mathbb{E}_{a\sim\pi(\cdot\mid s)}\bigl[Q_{h}(s,a)\bigr].$ (1)
@@ 3 Motif-based Continuous Dynamics (MCD) Discovery
Spectral decomposition has been widely studied in RL representation learning (Ren et al., 2022; Zhang et al., 2022; Shribak et al., 2024). We adopt this approach here to define motifs given the transition kernel and define rewards accordingly. Since spectral decomposition derives latents directly from the transition kernel without model assumptions, motif learning is thus independent of model assumptions and faithfully reflects the motifs present in the behavior data.
Given the motif definition, substituting Eq. 2 and the linearized reward model into Bellman equation (Eq. 2), we get:
$Q(s,a)$ $=r(s,a)+\gamma\!\!\int\!V(s^{\prime})P(s^{\prime}|s,a)\mathrm{d}s^{\prime}$
$=\phi(s,a)^{\top}\!\!\left[w+\gamma\!\!\int\!V(s^{\prime})\mu(s^{\prime})q(s^{\prime})\mathrm{d}s^{\prime}\right]$
$=\phi(s,a)^{\top}u,\!$ (3)
@@
$\pi(a|s)$ $=\arg\max_{\pi}\;\;\left[\mathbb{E}_{\pi}[Q(s,a)]+H(\pi)\right]$
$=\frac{\exp(\phi(s,a)^{\top}u)}{\sum_{a^{\prime}\in\mathcal{A}}\exp(\phi(s,a^{\prime})^{\top}u)},$ (4)
where the entropy $H(\pi):=-\sum_{a\in\mathcal{A}}\pi(a|s)\,\log\bigl(\pi(a|s)\bigr)$.
Proposition 3.2.
@@
From Def. 3.1 and Prop. 3.2, we conclude that the learned motifs and policies do not rely on model assumptions, yet the policies faithfully capture behavioral dynamics as action tendencies conditioned on state. Thus, our method is assumption-free while capturing dynamics, unlike classification/clustering methods (no dynamics) or dynamics-based methods (restrictive assumptions).
Next, we introduce how to learn $\phi(s,a)$ and $\mu(s^{\prime})$ (motif discovery), as well as $u$ (motif-based policy learning) from demonstrations. The learning procedure differs depending on the nature of the behavior data (i.e. discrete or continuous).
@@ 3.1 Discrete Version
Motif discovery. For discrete state-action spaces, there are two ways to learn the representation: (1) spectral methods such as singular value decomposition (SVD) (Golub and Reinsch, 1971; Golub and Van Loan, 2013; Trefethen and Bau, 2022); (2) spectral decomposition representation (Ren et al., 2022; HaoChen et al., 2021) with the objective $\phi(s,a),\mu(s^{\prime})=\arg\min_{\phi,\mu}||P(s^{\prime}|s,a)-\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime})||^{2}$. Here we choose the second method and set the auxiliary distribution as $q(s^{\prime})=\rho(s^{\prime})=\int\tau^{e}(s,a)da$, i.e. the marginalized state distribution. The resulting motif set $\phi(s,a)$ is then used in the subsequent policy learning stage.
@@
$\max_{u}\;\;\mathbb{E}_{(s,a)\sim\tau^{e}}\left[\log\pi(a|s)\right]$
$=$ $\max_{u}\;\;\mathbb{E}_{(s,a)\sim\tau^{e}}\left[\log\frac{\exp(\phi(s,a)^{\top}u)}{\sum_{a^{\prime}\in\mathcal{A}}\exp\left(\phi(s,a^{\prime})^{\top}u\right)}\right].$ (5)
@@ 3.2 Continuous Version
While learning from discrete data is relatively straightforward, the continuous case presents additional challenges for two main reasons. First, in the motif discovery step (Eq. 2), the decomposition $P(s^{\prime}|s,a)=\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime})$ is too restrictive to capture the complexity of continuous behavioral dynamics, such as pose transitions in freely moving animals (Weinreb et al., 2024). For example, consider a common and biologically plausible behavioral model: $s^{\prime}=h(s,a)+\epsilon$, where $h$ is a dynamics function and $\epsilon$ is Gaussian noise. This additive structure, widely used in behavioral modeling, contrasts with the multiplicative form $\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime})$, suggesting that parameterizations preserving additive relationships between $\{s,a\}$ and $s^{\prime}$ are more suitable. Second, in motif-based policy learning, for discrete datasets with a small action space, the denominator (partition function) in Eq. 3.1 is easy to compute. But for continuous data, the action space is infinite, making it infeasible to enumerate all actions and integrate. In light of these challenges, we adopt an alternative approach to learn the motifs and policy in continuous state-action spaces.
@@
$P(s^{\prime}|s,a)$ $=q(s^{\prime})\exp\left(\psi(s,a)^{\top}\nu(s^{\prime})-\log Z(s,a)\right),$
$Z(s,a)$ $=\int q(s^{\prime})\exp(\psi(s,a)^{\top}\nu(s^{\prime}))ds^{\prime},$ (6)
@@
Proposition 3.3 (Connection to Motif Definition).
@@
Given the EBM model in Eq. 3.2, the transition kernel can be approximated by $P(s^{\prime}|s,a)\approx\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime})$, where $\phi(s,a)\in\mathbb{R}^{d}$ is an explicit function of $\psi(s,a)$ and the partition function $Z(s,a)$, and $\mu(s^{\prime})\in\mathbb{R}^{d}$ is a function of $\nu(s^{\prime})$.
Appendix D contains the full proof and derivation of $\phi$ and $\mu$ in terms of $\psi$, $\nu$, and $Z$.
@@
To learn $\psi$ and $\nu$, we employ noise-contrastive estimation (NCE) (Wu et al., 2018; Ma and Collins, 2018; Gutmann and Hyvärinen, 2010, 2012), which enables optimization of unnormalized statistical models without explicitly computing the partition function. In this way, we sidestep the intractable computation of $Z(s,a)$ in Eq. 3.2 by solving
@@
$\max_{\psi,\,\nu}\;\;\mathbb{E}_{\begin{subarray}{c}(s,a)\sim\tau^{e},\\ s^{\prime}\sim P(\cdot|s,a),\\ s^{\prime\prime}_{i}\sim\rho\end{subarray}}\left[\log\frac{F(s,a,s^{\prime})}{F(s,a,s^{\prime})+\Sigma_{i=1}^{k}F(s,a,s^{\prime\prime}_{i}))}\right],$ (7)
@@
where $F(s,a,s^{\prime})=\exp(\psi(s,a)^{\top}\nu(s^{\prime}))$, $s^{\prime}$ denotes a positive sample drawn from the transition distribution $P(\cdot|s,a)$, and $s^{\prime\prime}_{i}$ for $i=1,\ldots,k$ are negative samples from marginalized state distribution $\rho(s)=\int\tau^{e}(s,a)da$.
@@
Connection to behavioral dynamics. With simple algebra, we obtain the quadratic potential function $P(s^{\prime}|s,a)\propto q(s^{\prime})\cdot\exp\left(\|\psi(s,a)\|^{2}/2\right)\cdot\exp\left(-\|\psi(s,a)-\nu(s^{\prime})\|^{2}/2\right)\cdot\left(\|\nu(s^{\prime})\|^{2}/2\right)$ from Eq. 3.2. By enforcing unit-norm constraints $\|\psi(s,a)\|^{2}=\|\nu(s^{\prime})\|^{2}=1$, assuming $Z(s,a)$ as a constant and $q(s^{\prime})$ as uniform distribution, as well as taking $\nu$ to be the identity map, we obtain a generalized Gaussian form: $P(s^{\prime}|s,a)=\frac{1}{Z}\exp\left(-\|s^{\prime}-\psi(s,a)\|^{2}\right)$, which aligns with commonly adopted assumptions in animal behavior modeling discussed earlier. Thus, Eq. 3.2 offers a more general framework that extends traditional dynamics models for studying behavior. Moreover, in practice, using Eq. 3.2 results in a unimodal distribution over $s^{\prime}$, which closely matches the empirical structure observed in $P(s^{\prime}|s,a)$ from animal behavior data. Thus, while Eq. 2 is theoretically valid in continuous domains, we adopt the parameterization in Eq. 3.2 for continuous state and action spaces, as it more effectively supports the learning of motif representations underlying animal behavior.
@@
Motif-based policy learning. After we obtain the representation $\psi(s,a)$ and $\nu(s^{\prime})$ from Eq. 7, we could theoretically get motif sets $\phi(s,a)$ expressed as the function of $\psi(s,a)$ and a normalizing term $Z(s,a)$ (see Appendix D). However, since in practice $Z(s,a)$ remains intractable, even with the optimal $\psi$ it is still hard to obtain $\phi$ exactly. Therefore, we introduce a mapping $f:\psi\rightarrow\phi$, parameterized with a neural network, and learn it via policy learning. Our aim is to learn a function $f$ so that $\phi=f(\psi)$ yields optimal basis functions of policy that best account for the animal behavior data. By applying $\phi=f(\psi)$ to Eq. 3, we obtain $Q(s,a)=f\bigl(\psi(s,a)\bigr)^{\top}u$.
@@
As mentioned earlier, learning both $f$ and $u$ using the MLE objective in Eq. 3.1 becomes intractable for continuous data, as the denominator involves integration over an unbounded continuous action space. This brings us back to the challenge of estimating an unnormalized energy function, $\pi(a|s)\propto\exp(Q(s,a))\propto\exp(f(\psi(s,a))^{\top}u)$. Thus, it’s reasonable to apply NCE here again,
@@
$\max_{f,\,u}\;\;\mathbb{E}_{\begin{subarray}{c}(s,a)\sim\tau^{e},\\ (s_{i}^{\prime},a^{\prime}_{i})\sim\tau^{e}\end{subarray}}\left[\log\frac{G(s,a)}{G(s,a)+\Sigma_{i=1}^{k}G(s,a^{\prime}_{i})}\right],$ (8)
@@
where $G(s,a)=\exp(f(\psi(s,a))^{\top}u)$, $\{s,a\}$ are positive samples and $\{s^{\prime}_{i},a^{\prime}_{i}\}$ are negative samples. Identifiability discussions are provided in Appendix J.
3.3 Understanding Animal Behavior via Motifs